### v1

# Cooperative Sheaf Neural Networks

Abstract
Sheaf diffusion has recently emerged as a promising design pattern for graph representation learning due to its inherent ability to handle heterophilic data and avoid oversmoothing. Meanwhile, cooperative message passing has also been proposed as a way to enhance the flexibility of information diffusion by allowing nodes to independently choose whether to propagate/gather information from/to neighbors. A natural question ensues: is sheaf diffusion capable of exhibiting this cooperative behavior? Here, we provide a negative answer to this question. In particular, we show that existing sheaf diffusion methods fail to achieve cooperative behavior due to the lack of message directionality. To circumvent this limitation, we introduce the notion of cellular sheaves over directed graphs and characterize their in- and out-degree Laplacians. We leverage our construction to propose Cooperative Sheaf Neural Networks (CSNNs). Theoretically, we characterize the receptive field of CSNN and show it allows nodes to selectively attend (listen) to arbitrarily far nodes while ignoring all others in their path, potentially mitigating oversquashing. Our experiments show that CSNN presents overall better performance compared to prior art on sheaf diffusion as well as cooperative graph neural networks.
1 Introduction
Graph neural networks (GNNs) have become the standard models for an array of sensitive machine learning over networked data, with far-reaching applications in, e.g., physics simulation (Sanchez-Gonzalez et al., 2020), recommender systems (Ying et al., 2018), and molecular modeling (Duvenaud et al., 2015; Gilmer et al., 2017). Most GNNs follow a message-passing paradigm, in which node representations are updated by aggregating features received from neighboring nodes. This approach assumes that every node participates in communication by both sending and receiving information.
A recent research direction focuses on generalizing GNNs and its message-passing paradigm through Sheaf Neural Networks (SNNs), which incorporate additional structure by modeling interactions using cellular sheaves. A cellular sheaf ${\mathcal{F}}$ over an undirected graph associates (i) vector spaces ${\mathcal{F}}(i)$ and ${\mathcal{F}}(e)$, known as stalks, to each vertex $i$ and each edge $e$, and (ii) a linear map $\mathcal{F}_{i\unlhd e}$, known as a restriction map, to each incident vertex-edge pair $i\unlhd e$. These mathematical constructs induce a continuous-time diffusion process over the graph, which we can discretize to derive neural architectures for relational data. Notably, Neural Sheaf Diffusion (NSD) (Bodnar et al., 2022) model learns these restriction maps and are naturally capable of avoiding oversmoothing while handling graphs with heterotrophic characteristics. In this context, the sheaf Laplacian serves as the main operator to propagate information, extending classical graph convolutional networks (Kipf and Welling, 2017).
In parallel, Cooperative GNNs (Finkelshtein et al., 2024) have been proposed as a way to increase the expressiveness of message-passing by relaxing the assumption that nodes must always send or receive information. In this framework, nodes can choose to ignore the features of their neighbors and/or choose not to transmit their own features. This selective communication allows the model to focus on long-range dependencies and mitigate over-squashing by increasing the sensitivity of node features to information at distant nodes.
Motivated by these advances, we propose to combine the strengths of both paradigms through Cooperative Sheaf Neural Networks (CSNNs). CSNNs extend SNNs by incorporating the cooperative message-passing mechanism in the context of sheaves. To do so, we enrich undirected graphs into directed ones, define the corresponding notion of cellular sheaves, and introduce the first formulation of a sheaf Laplacian for directed graphs in machine learning. While this doubling of graph edges increases the number of restriction maps to be learned, potentially leading to overfitting, we address this by learning two restriction maps per node, in a similar fashion to Bamberger et al. (2024), instead of full sheaves. In our formulation, we assign conformal maps ${\mathbf{S}}_{i}$ and ${\mathbf{T}}_{i}$ to each node $i$ — corresponding to their role as a source and target in an edge, respectively — rather than learning separate restriction maps for each incident edge $i\unlhd e$. We define that a node $i$ does not broadcast to its neighbors when ${\mathbf{S}}_{i}=0$, and does not listen to its neighbors when ${\mathbf{T}}_{i}=0$. These maps are learned independently at each layer of the CSNN architecture.
Our theoretical and empirical results demonstrate the effectiveness of CSNNs in alleviating over-squashing, while leveraging the expressiveness and relational richness of both sheaf-based and cooperative frameworks. Our experiments show CSNN performs better than the prior sheaf-based models and cooperative GNNs in standard node classification benchmarks.
In summary, our contributions are:
  1. 1.
    We provide a novel sheaf Laplacian over directed graphs, expanding the applications of sheaf theory in deep learning from undirected GNNs to directed GNNs.
  2. 2.
    For each layer $t$ in our CSNN model, each node is affected by information from nodes at distance up to $2t$-hop neighbors (Proposition 4.2), instead of up to $t$-hop neighbors in usual GNNs. Moreover, we can learn a sheaf in which the feature of a node $i$ at layer $t$ is highly sensitive to the initial feature of a node $j$, where $t$ is also the distance between $i$ and $j$ (Proposition 4.3).
  3. 3.
    We present a new SNN model that preserves its characteristics of avoiding oversmoothing and better perform on heterotrophic graphs, and additionally, mitigates overfitting and over-squashing, by using a similar approach proposed by Bamberger et al. (2024) but with an architecture closer to the NSD presented in Bodnar et al. (2022).
2 Background
We begin by introducing the core concepts of cellular sheaves over undirected graphs.
Definition 2.1.
A cellular sheaf $(G,\mathcal{F})$ over a (undirected) graph $G=(V,E)$ associates:
  1. 1.
    Vector spaces $\mathcal{F}(i)$ to each vertex $i\in V$ and $\mathcal{F}(e)$ to each edge $e\in E$, called stalks.
  2. 2.
    Linear maps $\mathcal{F}_{i\unlhd e}\!:\!\mathcal{F}(i)\to\mathcal{F}(e)$ to each incident vertex-edge pair $i\unlhd e$, called restriction maps.
Hereafter, we assume all vertex and edge stalks are isomorphic to ${\mathbb{R}}^{d}$. If all restriction maps are equal to the identity map, we say the cellular sheaf is constant. Furthermore, if $d\!=\!1$, the sheaf is said trivial.
Hansen and Ghrist (2021) interpret the node stalks ${\mathcal{F}}(i)$ as the space of private opinions of a person $i$. If $e$ is an edges between $i$ and $j$, then ${\mathcal{F}}(e)$ represents the space of public opinions discussed between these two nodes. We will leverage this allegory to provide intuitive accounts of technical definitions.
Given the importance of the Laplacian operator for graph representation learning, it is instrumental to define the Laplacian for undirected cellular sheaves – a key concept in the design of SNNs. Towards this end, the definition below introduces the domain and codomain of the coboundary operator.
Definition 2.2.
The space of 0-cochains, denoted by $C^{0}(G,{\mathcal{F}})$, and the space of 1-cochains, $C^{1}(G,{\mathcal{F}})$, of a cellular sheaf $(G,{\mathcal{F}})$ are given by
$C^{0}(G,{\mathcal{F}})=\bigoplus_{i\in V}{\mathcal{F}}(i)\mbox{ and }C^{1}(G,{\mathcal{F}})=\bigoplus_{e\in E}{\mathcal{F}}(e).$
where $\oplus$ denotes the (external) direct sum.
Now, for each $e\in E$ choose an orientation $e\!=\!i\!\to\!j$ and consider the coboundary operator $\delta\!:\!C^{0}(G,{\mathcal{F}})\to C^{1}(G,{\mathcal{F}})$ defined by $(\delta\mathbf{X})_{e}=\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}-\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}$. Then, the sheaf Laplacian is defined by $L_{{\mathcal{F}}}=\delta^{\top}\delta$. If $\mathcal{F}$ is the trivial sheaf, $\delta^{\top}$ can be seen as the incidence matrix, recovering the $n\times n$ graph Laplacian. A more explicit way to describe the Laplacian is the following:
Definition 2.3.
The sheaf Laplacian of a cellular sheaf $(G,\mathcal{F})$ is the linear operator $L_{\mathcal{F}}:C^{0}(G,\mathcal{F})\to C^{0}(G,\mathcal{F})$ that, for a 0-cochain $\mathbf{X}\in C^{0}(G,\mathcal{F})$, outputs
$L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\mathcal{F}_{i\unlhd e}^{\top}\left(\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}-\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\right)\qquad\forall i\in V.$ (1)
The Laplacian $L_{\mathcal{F}}$ can also be seen as a positive semidefinite matrix with diagonal blocks $L_{ii}=\sum_{i\unlhd e}\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{i\unlhd e}$ and non-diagonal blocks $L_{ij}=L_{ij}^{\top}=-\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{i\unlhd e}$.
Note that $\ker L_{{\mathcal{F}}}=\{\mathbf{X}\in C^{0}(G,{\mathcal{F}})\>|\>\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\,\forall e=\{i,j\}\in E(G)\}$. Using the opinions analogy, this can be understood as the space of public agreement between all pairs of neighboring nodes $i$ and $j$. Note that $i$ and $j$ can have distinct opinions about the same topic on their respective private opinion spaces $\mathcal{F}(i)$ and $\mathcal{F}(j)$; however, when they publicly discuss about this topic, they may prefer to not manifest their true opinion. This apparent consensus lies in $\operatorname{Ker}L_{{\mathcal{F}}}$.
Vector bundles. When restriction maps are orthogonal, we call the sheaf a vector bundle. In this case, $L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\left(\mathbf{x}_{i}-\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\right)$, for each vertex $i\in V$. Bodnar et al. (2022) found vector bundles lead to computationally more efficient NSD models and superior empirical performance.
Flat vector bundles. To further enhance the scalability of sheaf-based models, Bamberger et al. (2024) considers the special case of flat vector bundles in which to every node $i$ is assigned an orthogonal map ${\mathbf{O}}_{i}$, which replaces $\mathcal{F}_{i\unlhd e}$ for any adjacent edge $e$. This is a clear computational advantage since we just need to learn the orthogonal maps ${\mathbf{O}}_{i}$ instead of all restriction maps. In this case, $L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\left(\mathbf{x}_{i}-{\mathbf{O}}_{i}^{\top}{\mathbf{O}}_{j}\mathbf{x}_{j}\right)$, for each vertex $i\in V$.
3 Cellular sheaves for directed graphs
In this section, we construct a sheaf Laplacian for directed graphs. To the best of our knowledge, we are the first to discuss this in the machine learning community. Moreover, we have found a single paper proposing a (symmetric) sheaf Laplacian in a similar context (Sumray et al., 2024). Before deepening the discussion on the sheaf Laplacian, we introduce the definition of sheaves over directed graphs.
Cellular sheaves over directed graphs must distinguish the restriction map where $i$ is the source node of an edge from the restriction map where $i$ is the target node of an edge. Therefore, we change the edge notation from $e$ to $ij$ and $ji$ to make this distinction explicit. See Figure 1 for an illustration.
Definition 3.1.
A cellular sheaf $(G,\mathcal{F})$ over a (directed) graph $G=(V,E)$ associates:
  1. 1.
    Vector spaces $\mathcal{F}(i)$ to each vertex $i\in V$ and $\mathcal{F}(ij)$ to each edge $ij\in E$, called stalks.
  2. 2.
    A linear map $\mathcal{F}_{i\unlhd ij}:\mathcal{F}(i)\to\mathcal{F}(ij)$ for each incident vertex-edge pair $i\unlhd ij$ and a linear map $\mathcal{F}_{i\unlhd ji}:\mathcal{F}(i)\to\mathcal{F}(ji)$ for each incident vertex-edge pair $i\unlhd ji$, called restriction maps.
Figure 1: On the left we have a cellular sheaf shown for a single edge of an undirected graph with stalks isomorphic to $\mathbb{R}^{2}$. The restriction maps $\mathcal{F}_{i\unlhd e},\mathcal{F}_{j\unlhd e}$ move the vector features between these spaces. On the right we have the analogous situation for a sheaf on a single pair of directed edges. Then there are four, possibly distinct, restriction maps $\mathcal{F}_{i\unlhd ij},\mathcal{F}_{i\unlhd ji},\mathcal{F}_{j\unlhd ij},\mathcal{F}_{j\unlhd ji}$.
Again, a common dimension $d$ is fixed for all node and edge stalks. Under the opinion analogy, directed sheaves allow for the restriction maps $\mathcal{F}_{i\unlhd ij}$ and $\mathcal{F}_{i\unlhd ji}$ to be different: we can have both edges $ij$ and $ji$ meaning that both people know each other but $j$ may not listen to $i$ $(\mathcal{F}_{i\unlhd ij}=0)$ or be less influenced by $i$ than $i$ is influenced by $j$, for instance.
For directed graphs, it is common to define both in- and out-degree Laplacians (Agaev and Chebotarev, 2005). Given a directed graph with possibly asymmetric adjacency matrix $A$, the out-degree Laplacian is $L^{\text{out}}:=D^{\text{out}}-A$ and the in-degree Laplacian is $L^{\text{in}}:=D^{\text{in}}-A$, with $D^{\text{in}}$ and $D^{\text{out}}$ denoting the diagonal matrices containing in- and out-degree of nodes in $V(G)$. Definition 3.2 generalizes these notions to directed sheaves. Note that if $(G,{\mathcal{F}})$ is the trivial sheaf, then $L^{\text{out}}_{\mathcal{F}}=(L^{\text{out}})^{\top}$ and $L^{\text{in}}_{\mathcal{F}}=L^{\text{in}}$.
Definition 3.2.
The out-degree sheaf Laplacian of a cellular sheaf $(G,\mathcal{F})$ is the linear operator $L^{\text{out}}_{\mathcal{F}}:C^{0}(G,\mathcal{F})\to C^{0}(G,\mathcal{F})$ that, for a 0-cochain $\mathbf{X}\in C^{0}(G,\mathcal{F})$, outputs
$L^{\text{out}}_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{j\in N(i)}\left(\mathcal{F}_{i\unlhd ij}^{\top}\mathcal{F}_{i\unlhd ij}\mathbf{x}_{i}-\mathcal{F}_{i\unlhd ji}^{\top}\mathcal{F}_{j\unlhd ji}\mathbf{x}_{j}\right),\qquad\forall i\in V.$ (2)
The in-degree sheaf Laplacian of a cellular sheaf $(G,\mathcal{F})$ is the linear operator $L^{\text{in}}_{\mathcal{F}}:C^{0}(G,\mathcal{F})\to C^{0}(G,\mathcal{F})$ that, for a 0-cochain $\mathbf{X}\in C^{0}(G,\mathcal{F})$, outputs
$L^{\text{in}}_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{j\in N(i)}\left(\mathcal{F}_{i\unlhd ji}^{\top}\mathcal{F}_{i\unlhd ji}\mathbf{x}_{i}-\mathcal{F}_{i\unlhd ij}^{\top}\mathcal{F}_{j\unlhd ij}\mathbf{x}_{j}\right),\qquad\forall i\in V.$ (3)
We anticipate that working directly with sheaves will lead to scalability issues and overfitting since a map would be learned for each node-edge pair, as reported in Bodnar et al. (2022) for undirected sheaf models. Bamberger et al. (2024) alleviate this issue using flat bundles: we only have to learn one orthogonal map to each node since ${\mathbf{O}}_{i}=\mathcal{F}_{i\unlhd e}$ for every adjacent $e$. Since the graphs are directed, we need to distinguish if the adjacent edge is pointing to $i$ or leaving $i$, then for each node-edge pair we learn two maps. More precisely, for each node $i$, we have the same source conformal map ${\mathbf{S}}_{i}=\mathcal{F}_{i\unlhd ij}$ and a target conformal map ${\mathbf{T}}_{i}=\mathcal{F}_{i\unlhd ji}$, for all neighbors $j$ of $i$.
4 Cooperative Sheaf Neural Networks
In this section, we leverage the sheaf Laplacians in Definition 3.2 to introduce Cooperative Sheaf Neural Networks (CSNNs). To exploit the asymmetric communication induced by directed sheaves, we first convert our input (undirected) graph into a directed one in the following way: every time there is an edge between nodes $i$ and $j$, an arbitrary direction is chosen to form the graph $G$, then we also consider the reversed graph $\overline{G}$. Finally, we use a sheaf $(G\cup\overline{G},{\mathcal{F}})$ whose restriction maps $\mathcal{F}_{i\unlhd ij}$ and $\mathcal{F}_{i\unlhd ji}$ control the information from $i$ and to $i$, respectively.
Our goal is to create a sheaf diffusion model capable of exhibiting cooperative behavior. In other words, nodes should be able to decide how they participate in message passing, choosing whether to broadcast their information or to listen from the neighbors. For undirected sheaf models, it is easy to see that this behavior is unachievable since both the message from $i$ to $j$ through an undirected edge $e$ and its converse depend directly on $\mathcal{F}_{i\unlhd e}^{\top}$ — and setting $\mathcal{F}_{i\unlhd e}^{\top}$ to zero cancels out messages in both directions (c.f., Equation 1).
We design CSNN to achieve this flexible behavior by composing the out-degree and the transposed in-degree sheaf Laplacians. For each node $i$, we will define a source conformal map ${\mathbf{S}}_{i}$ and a target conformal map ${\mathbf{T}}_{i}$, for all neighbor $j$ of $i$. In this case, out-degree sheaf Laplacian simplifies to
$L^{\text{out}}_{\mathcal{F}}(\mathbf{X})_{i}$ $:=\textstyle\sum_{j\in N(i)}\left({\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}\mathbf{x}_{i}-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\textbf{x}_{j}\right),$ (4)
while the transpose of the in-degree sheaf Laplacian is
$((L^{\text{in}}_{\mathcal{F}})^{\top}(\mathbf{X}))_{i}:=\textstyle\sum_{j\in N(i)}\left({\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}\mathbf{x}_{i}-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\mathbf{x}_{j}\right).$ (5)
Note these matrices have a block structure, with diagonals $({L^{\text{in}}_{\mathcal{F}}})^{\top}_{ii}=\sum{\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}$, $({L^{\text{out}}_{\mathcal{F}}})_{ii}=\sum{\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}$, and remaining blocks $(L^{\text{in}}_{{\mathcal{F}}})^{\top}_{ij}=(L^{\text{out}}_{{\mathcal{F}}})_{ij}=-{\mathbf{T}}^{\top}_{i}{\mathbf{S}}_{j}$. In practice, we use their normalized versions
$\Delta_{{\mathcal{F}}}^{\text{out}}=D_{\text{out}}^{-\frac{1}{2}}L_{{\mathcal{F}}}^{\text{out}}D_{\text{out}}^{-\frac{1}{2}}\quad\text{ and }\quad(\Delta_{{\mathcal{F}}}^{\text{in}})^{\top}=D_{\text{in}}^{-\frac{1}{2}}(L_{{\mathcal{F}}}^{\text{in}})^{\top}D_{\text{in}}^{-\frac{1}{2}},$
where $D_{\text{in}},D_{\text{out}}$ are the block-diagonals of the in and out Laplacians, respectively. We point out that, since conformal maps are of the form ${\mathbf{S}}_{i}=C_{{\mathbf{S}}_{i}}Q_{i}$ and ${\mathbf{T}}_{i}=C_{{\mathbf{T}}_{i}}R_{i}$, for some orthogonal matrices $Q_{i}$, $R_{i}$ and scalars $C_{{\mathbf{S}}_{i}},\>C_{{\mathbf{T}}_{i}}$, the block diagonals are just scalars times an identity matrix, hence the normalization is both numerically stable and computationally efficient.
The CSNN model is an adaptation of the NSD model from Bodnar et al. (2022), which is given by
$\mathbf{X}_{t+1}=(1+\varepsilon)\mathbf{X}_{t}-\sigma(\Delta_{{\mathcal{F}}(t)}(\mathbf{I}\otimes\mathbf{W}_{1,t})\mathbf{X}_{t}\mathbf{W}_{2,t}),$
where $\Delta_{{\mathcal{F}}(t)}$ is the sheaf Laplacian Equation 1 for undirected graphs. Concisely, we define a CSNN layer from the NSD model substituting the sheaf Laplacian of the undirected graph by the composition of normalized in- and out-Laplacians:
$\textbf{X}_{t+1}=(1+\varepsilon)\textbf{X}_{t}-\sigma((\Delta_{{\mathcal{F}}(t)}^{in})^{\top}\Delta_{{\mathcal{F}}(t)}^{out}(\textbf{I}\otimes\textbf{W}_{1,t})\textbf{X}_{t}\textbf{W}_{2,t}),$ (6)
where $\varepsilon\in[-1,1]^{nd}$, $\sigma$ is an activation, and ${\mathbf{W}}_{1}\in{\mathbb{R}}^{d\times d},{\mathbf{W}}_{2}\in{\mathbb{R}}^{h\times h}$ are weight matrices. We provide a pseudo-code on how we implement the CSNN model in Algorithm 1.
Algorithm 1 Cooperative Sheaf Diffusion
Input: Node features $X_{0}\in{\mathbb{R}}^{n\times s}$, graph $G$, number of layers $L$.
Output: Class probabilities for each node $X_{\mathrm{final}}$
1:$\mathbf{X}_{0}=\sigma(X_{0}W_{0}+b_{0})$triangleright $\mathbf{X}_{0}\in{\mathbb{R}}^{nd\times h}$
2:for $\ell=0,\dots,L-1$ do
3:     ${\mathbf{S}}_{i}=\eta(\mathbf{X}_{\ell},G,i),\>\forall i\in V$triangleright Learn source maps
4:     ${\mathbf{T}}_{i}=\phi(\mathbf{X}_{\ell},G,i),\>\forall i\in V$triangleright Learn target maps
5:     $L_{{\mathcal{F}}}^{\text{in}},L_{{\mathcal{F}}}^{\text{out}}=\mathrm{LaplacianBuilder}(\{{\mathbf{S}}_{i}\}_{i=1}^{n},\{{\mathbf{T}}_{i}\}_{i=1}^{n})$
6:     $\mathbf{X}_{\ell+1}=\mathrm{Diffusion}(\mathbf{X}_{\ell},L_{{\mathcal{F}}}^{\text{in}},L_{{\mathcal{F}}}^{\text{out}})$triangleright Equation 6
7:end for
8:Return: $\mathrm{READOUT}(\mathbf{X}_{L})$
The first step of each layer $\ell$ consists in computing the conformal maps ${\mathbf{T}}_{i,\ell}$ and ${\mathbf{S}}_{i,\ell}$. We do this through learnable functions ${\mathbf{S}}_{i,\ell}=\eta(G,\mathbf{X}_{\ell},i)$ and ${\mathbf{T}}_{i,\ell}=\phi(G,\mathbf{X}_{\ell},i)$, $\forall i\in V$. As in Bamberger et al. (2024), we can use any neural network to learn the restriction maps. In detail, we use Householder reflections (Mhammedi et al., 2017; Obukhov, 2021) to compute orthogonal maps and multiply them by a learned positive constant for each node. Finally, the Laplacians $(\Delta_{{\mathcal{F}}}^{in})^{\top}$ and $\Delta_{{\mathcal{F}}}^{out}$ are computed. In light of Proposition 4.1, it is clear that this modification was intended to create a model in which a node can ignore its neighbors or choose not to send them information.
In particular, CSNN allows each node $i$ to play one of the four following roles in information diffusion:
  • STANDARD: This is the standard message-passing paradigm. The node $i$ propagates information from and to its neighbors. Mathematically, this corresponds to ${\mathbf{S}}_{i},{\mathbf{T}}_{i}\neq 0$.
  • LISTEN: The node $i$ receives information from broadcasting neighbors, i.e. ${\mathbf{S}}_{i}=0$ and ${\mathbf{T}}_{i}\neq 0$.
  • BROADCAST: Node $i$ propagates information to the neighbors that listen, i.e. ${\mathbf{T}}_{i}=0$ and ${\mathbf{S}}_{i}\neq 0$.
  • ISOLATE: The node $i$ neither listens nor broadcasts, i.e. ${\mathbf{S}}_{i}=0={\mathbf{T}}_{i}$.
Figure 2: ${\mathbf{O}}_{i}=0$ creates the effect of isolating the node $i$. Directed edges provide the possibility of considering ${\mathbf{S}}_{i}=0$ and ${\mathbf{T}}_{i}=0$, performing listen and broadcast, separately.
We highlight the importance of considering the directions: for undirected graphs, we have a single map ${\mathbf{O}}_{i}$ for each node $i$, where ${\mathbf{O}}_{i}=0$ could only mean that $i$ does not communicate at all. In other words, the possible actions are only standard and isolate. We illustrate this in Figure 2.
Goal. Ideally, our model must satisfy the following: if $i$ does not listen, then its update cannot depend of $\mathbf{x}_{j}$, $\forall j\in V,j\neq i$. If $i$ has a non-broadcasting neighbor $k$, then its update cannot depend on $\mathbf{x}_{k}$. The composition of out- and (transposed) in-Laplacians has this property, as we prove in Proposition 4.1.
Proposition 4.1.
If the target map of $i$ is zero, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}=0$. If ${\mathbf{S}}_{k}=0$ for some neighbor $k$ of $i$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}$ does not depend on $\mathbf{x}_{k}$.
Proof.
We have that $(L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}$ valued at a given vertex $i$ is:
$\begin{split}\sum_{j\in N(i)}\left({\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}\left(\sum_{j\in N(i)}\left({\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}\mathbf{x}_{i}-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\mathbf{x}_{j}\right)\right)-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\left(\sum_{u\in N(j)}\left({\mathbf{S}}_{j}^{\top}{\mathbf{S}}_{j}\mathbf{x}_{j}-{\mathbf{T}}_{j}^{\top}{\mathbf{S}}_{u}\mathbf{x}_{u}\right)\right)\right)\end{split}$ (7)
So ${\mathbf{T}}_{i}=0$ (i.e. $i$ does not listen) implies $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}=0$. If $i$ listens, but a certain neighbor $k$ does not broadcast, i.e., ${\mathbf{T}}_{k}=0$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}$ is
$\begin{split}\sum_{j\in N(i)\setminus k}\left({\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}\left(\sum_{j\in N(i)\setminus k}\left({\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}\mathbf{x}_{i}-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\mathbf{x}_{j}\right)\right)-{\mathbf{T}}_{i}^{\top}{\mathbf{S}}_{j}\left(\sum_{u\in N(j)\setminus k}\left({\mathbf{S}}_{j}^{\top}{\mathbf{S}}_{j}\mathbf{x}_{j}-{\mathbf{T}}_{j}^{\top}{\mathbf{S}}_{u}\mathbf{x}_{u}\right)\right)\right)\end{split}$
Since the sum does not go through the index $k$, $\mathbf{x}_{k}$ is not a component in $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}$. ∎
Moreover, as a consequence of Equation 7, our model has the ability to reach longer distances:
Proposition 4.2.
In each layer $t$, the features of a node can be affected by features of nodes up to $2t$-hop neighbors.
Proof.
Let $t=1$, and fix a node $i$. Then we are essentially just using the composition described in Equation 7 (up to normalization and learnable weights). In the equation we have a sum running over all neighbors $j$ of $i$ and another sum running over all neighbors $u$ of each $j$. So $u$ can be a $2$-hop neighbor of $i$ and we have that $i$ was updated with information from up to $2$-hop neighbors. Similarly, the node $u$ is updated by up to $2$-hop neighbors. Therefore, in the second layer $t=2$, $i$ was updated with information from up to $4$-hop neighbors.
If $t=n$, assume by induction that each node $i$ receives information from it $2n$-hop neighbors. In the next layer $n+1$, $i$ will be updated by its $n$-update of its $2$-hop neighbors. Let $j$ be a node in the $2$-hop neighborhood of $i$. By the inductive hypothesis, $j$ receives information from its $2n$-hop neighbors, whose distance to $i$ is up to $2n+2=2(n+1)$, concluding the proof by induction. ∎
Proposition 4.2 shows that CSNN can reach longer distances: in most GNNs if $t$ is the distance between two nodes $i$ and $j$, then they can only communicate after $t$ layers. CSNN enables communication between these nodes after $\lceil t/2\rceil$ layers. Moreover, the following result shows that CSNNs with $t$ layers are capable of making $i$ and $j$ communicate while ignoring all the other nodes on a path from $i$ to $j$, as proved in the next result.
Proposition 4.3.
Let $i$ and $j$ be nodes at a distance $t$. In CSNN, $i$ can learn to ignore all the $t-1$ nodes in the shortest path from $i$ to $j$ while receiving the information from $j$ in the $t$-layer. Moreover, if we choose a path with $n>t-1$ nodes between $i$ and $j$, then $i$ receives the information from $j$ in the $(n+1)$-layer.
Proof.
Choose a path from $i$ to $j$. So there are $t-1$ vertices between $i$ an $j$, say $v_{1},...,v_{t-1}$. In the first layer, let ${\mathbf{S}}_{j}$ and ${\mathbf{T}}_{v_{t-1}}$ be different of zero and all other source and target maps equal zero.
This results in $((\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out})_{k}=0$ for every $k\neq v_{t-1}$ and $((\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out})_{v_{t-1}}$ depends only of $x_{j}$. So, except for $x_{v_{t-1}}$, the values $x_{k}$ are not updated.
In the second layer, let ${\mathbf{S}}_{v_{t-1}}$ and ${\mathbf{T}}_{v_{t-2}}$ different of zero, and all other maps equal zero. This results in $((\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out})_{k}=0$ for every $k\neq v_{t-1}$, where $((\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out})_{v_{t-1}}$ depends only of the $x^{(1)}_{v_{t-1}}$ that was updated in the previous layer and depends only of $x_{j}$.
We continue this reasoning until the $t$-layer, in which we make ${\mathbf{S}}_{v_{1}}$ and ${\mathbf{T}}_{i}$ different of zero, and all other source maps equal zero. This results in $((\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out})_{k}=0$ for every $k\neq i$, where $((\Delta_{{\mathcal{F}}}^{in})^{\top}\Delta_{{\mathcal{F}}}^{out})_{v_{1}}$ depend only of the $x^{(t-1)}_{v_{1}}$, which going backwards depends only of the original $x_{j},$ up to transformations given by the target and source maps. ∎
The above shows that by learning the right sheaves in each layer, if a node $i$ wants to receive information only from node $j$ at distance $t$, then it can ignore all other nodes until it reaches the $t$-layer, where it will get the information from node $j$ that was updated only by itself, layer by layer. Intuitively, a node $i$ will be updated by a neighbor $k$ only if it is able do listen and $k$ is able to broadcast. This is the setup we use to go from one node to the next one without updating anything else until it reaches the final node. Although we need all the $t$-layers, this is an advantage compared to standard message-passing since we have the possibility to ignore other nodes.
Example 4.4.
Consider a graph formed by vertices $\{1,2,3,4,5\}$ and edges $\{12,23,34,45,54,43,32,21\}$. We follow the proof of 4.3 to check we can propagate a message from $5$ to $1$, while $1$ ignores all other nodes. In the first layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{4}$ and ${\mathbf{S}}_{5}$. A simple verification gives that $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 4$. So $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 4$. If $k=4$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{4}=-2{\mathbf{T}}_{4}^{\top}{\mathbf{S}}_{5}\textbf{x}_{5}^{(0)},$ with $\textbf{x}_{k}^{(t)}$ denoting the feature vector of $k$ at layer $t$, thus $\textbf{x}_{5}^{(0)}$ will be the only feature vector to influence $\textbf{x}_{4}^{(1)}$.
In the second layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{3}$ and ${\mathbf{S}}_{4}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 3$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{3}=-2{\mathbf{T}}_{3}^{\top}{\mathbf{S}}_{4}\textbf{x}_{4}^{(1)}$. Thus $\textbf{x}_{4}^{(1)}$ will be the only feature vector to influence $\textbf{x}_{3}^{(2)}$.
In the third layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{2}$ and ${\mathbf{S}}_{3}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 2$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{2}=-2{\mathbf{T}}_{2}^{\top}{\mathbf{S}}_{3}\textbf{x}_{3}^{(3)}$. Thus $\textbf{x}_{3}^{(2)}$ will be the only feature vector to influence $\textbf{x}_{2}^{(3)}$.
In the fourth and last layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{1}$ and ${\mathbf{S}}_{2}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 1$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{1}=-2{\mathbf{T}}_{1}^{\top}{\mathbf{S}}_{2}\textbf{x}_{2}^{(3)}$. Thus $\textbf{x}_{1}^{(4)}$ is influence only by $\textbf{x}_{2}^{(3)}$, which was influenced only by $\textbf{x}_{3}^{(2)}$, which in turn was influenced only by $\textbf{x}_{4}^{(1)}$, which receive information only from $\textbf{x}_{5}^{(0)}$. Consequently, $\textbf{x}_{1}^{(4)}$ is affected by $\textbf{x}_{5}^{(0)}$ while ignoring the feature vectors of all other nodes in all other layers.
Figure 3: Illustration of Example 4.4. At layer $\ell$, we consider that all maps but ${\mathbf{T}}_{5-\ell}$ and ${\mathbf{S}}_{5-(\ell-1)}$ are 00, enabling the flow of information from right to left following the bottom edges.
Observe that the derivative of $\textbf{x}_{1}^{(4)}$ in relation to $\textbf{x}_{5}^{(0)}$ can be as high as the values of the non-zero ${\mathbf{T}}_{i}$ and ${\mathbf{S}}_{i}$ permit. This shows our model can mitigate over-squashing, which refers to the failure of an information propagating to distance nodes. Di Giovanni et al. (2023) and Topping et al. (2022), studied over-squashing in message passing neural networks through a bound on the Jacobian
$\left|\frac{\partial\textbf{x}_{i}^{(t)}}{\partial\textbf{x}_{j}^{(0)}}\right|\leq c^{t}\hat{A}^{t}_{ij},$ (8)
where $t$ is the layer, $c$ is a constant that depends on the architecture of the model, and $\hat{A}$ the normalized adjacency matrix. Moreover, over-squashing occurs when we have a small derivative $\nicefrac{{\partial\textbf{x}_{i}^{(t)}}}{{\partial\textbf{x}_{j}^{(0)}}}$, since it means that after $t$ layers, the feature at $i$ is mostly insensitive to the information initially contained at $j$, i.e., the information was not propagated properly. Proposition 4.3 says that the feature at $i$ can be sensitive to the information initially contained at $j$, independently of the distance, if the right sheaves are learned and sufficient layers are used.
5 Experimental Results
This section provides real-world experiments to validate the proposed model. We use the recently proposed benchmarks from Platonov et al. (2023) to test our model capacity on dealing with node classification tasks focusing on heterophilic settings, which are notoriously challenging for GNNs. Additionally, we check the capacity of CSNN to deal with over-squashing on a synthetic experiment.
5.1 Over-squashing
In order to verify our theoretical results on the capacity of CSNN to alleviate over-squashing, we reproduce the NeighborsMatch problem proposed by Alon and Yahav (2020), using the same framework. The datasets consist of binary trees of depth $r$, with the root node as the target, the leaves containing its possible labels, and the leaf with the same number of neighbors as the target node containing its true label. We provide the parameters used for this task in Appendix B.
Figure 4 shows that GCN and GIN fail to fit the datasets starting from $r=4$ and GAT and GGNN fail to fit the datasets starting from $r=5$. These models suffered from over-squashing and were not able to distinguish between different training examples, while the CSNN model reached perfect accuracy for all tested $r$.
[…]
Figure 4: Accuracy across problem radius (tree depth) in the NeighborsMatch problem. CSNN is able to perfectly fit the training data, surpassing the bottleneck.
Alon and Yahav (2020) justified the difference performance between the four originally considered GNNs with the way the node features are updated: on one hand, GCN and GIN aggregate all neighbor information before combining it with the representation of the target node, forcing them to compress all incoming information into a single vector. On the other hand, GAT uses attention to selectively weigh messages based on the representation of the target node, allowing it to filter out irrelevant edges and only compress information from a subset of the neighbors. So models like GAT (and GGNN) that compress less information per step can handle higher $r$ better than GCN and GIN.
This experiment shows that the CSNN model is more efficient in ignoring irrelevant nodes and can avoid loosing relevant information. Moreover, Proposition 4.3 provides theoretical support for this result, as it states that the model can learn to listen only to the nodes along a path between distant nodes $i$ and $j$, enabling selective communication to diminish noise impact.
5.2 Node Classification
Datasets. We evaluate our model on the five heterophilic graphs from Platonov et al. (2023). The datasets include both binary (minesweeper, tolokers, questions) and multiclass (roman-empire, amazon-ratings) classification tasks. For binary classification datasets, we report ROC AUC, while for multiclass classification datasets we report accuracy.
Baselines. As baselines, we use GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2017) and GT (Shi et al., 2020), together with the variations GAT-sep and GT-sep, which concatenate the representation of a node to the mean of its neighbors, instead of summing them (Zhu et al., 2020). These are all classical baselines used in Platonov et al. (2023) to compare against GNN architectures specifically developed for heterophilic settings, and that achieve the best performance in most cases. We also compare CSNN against recent models such as CO-GNN (Finkelshtein et al., 2024), NSD (Bodnar et al., 2022), and BuNN (Bamberger et al., 2024).
Setup. For all datasets, we use the 10 fixed splits proposed by Platonov et al. (2023). For binary classification tasks, we report the mean and standard deviation of ROC AUC, and for multiclass classification tasks, we report mean and standard deviation of accuracy. Refer to Appendix A for further details.
Results. Table 1 shows that CSNN is consistently better than NSD in all tasks, while outperforms other models in all datasets but amazon-ratings. In general, our model achieves the best performance in 4 out of 5 datasets, with an exceptionally high score in minesweeper and a low score in amazon-ratings.
Table 1: Performance comparison on different datasets. Accuracy is reported for roman-empire and amazon-ratings, ROC AUC is reported for the remaining datasets.
Model roman-empire amazon-ratings minesweeper tolokers questions
GCN 73.69 ± 0.74 48.70 ± 0.63 89.75 ± 0.52 83.64 ± 0.67 76.09 ± 1.27
SAGE 85.74 ± 0.67 53.63 ± 0.39 93.51 ± 0.57 82.43 ± 0.44 76.44 ± 0.62
GAT 80.87 ± 0.30 49.09 ± 0.63 92.01 ± 0.68 83.70 ± 0.47 77.43 ± 1.20
GAT-sep 88.75 ± 0.41 52.70 ± 0.62 93.91 ± 0.35 83.78 ± 0.43 76.79 ± 0.71
GT 86.51 ± 0.73 51.17 ± 0.66 91.85 ± 0.76 83.23 ± 0.64 77.95 ± 0.68
GT-sep 87.32 ± 0.39 52.18 ± 0.80 92.29 ± 0.47 82.52 ± 0.92 78.05 ± 0.93
CO-GNN 89.44 ± 0.50 54.20 ± 0.34 97.35 ± 0.63 84.84 ± 0.96 75.97 ± 0.89
NSD 80.41 ± 0.72 42.76 ± 0.54 92.15 ± 0.84 78.83 ± 0.76 69.69 ± 1.46
BuNN 91.75 ± 0.39 53.74 ± 0.51 98.99 ± 0.16 84.78 ± 0.80 78.75 ± 1.09
CSNN 92.63 ± 0.50 52.07 ± 1.00 99.07 ± 0.25 85.45 ± 0.53 79.31 ± 1.22
5.3 Ablation: How important are the actions learned for the nodes
One can argue that cellular sheaves with orthogonal maps, considering oriented edges, may be sufficient to modulate message passing in heterophilic settings and learn meaningful node representations. To assess whether the actions learned through conformal maps in CSNN are necessary for strong performance, or if a simpler approach using only source and target bundles suffices, a model we call Directed Sheaf Network (DSN), we conduct an ablation study on heterophilic node classification tasks.
Setup. We compare CSNN, which learns conformal maps, to DSN, which only learns source and target bundles, on the five heterophilic datasets using the same splits and evaluation metrics as before.
Results. As shown in Table 2, CSNN outperforms DSN on four out of five datasets, with DSN showing a marginal advantage only on minesweeper. This indicates that learning conformal maps is generally beneficial for node classification in heterophilic graphs.
Table 2: Comparison between CSNN and DSN. We can see that in most cases, it is indeed useful to learn sheaves with conformal maps instead of bundles.
Model roman-empire amazon-ratings minesweeper tolokers questions
CSNN 92.63 ± 0.50 52.07 ± 1.00 99.07 ± 0.25 85.45 ± 0.53 79.31 ± 1.22
DSN 92.40 ± 0.52 51.19 ± 0.54 99.20 ± 0.26 84.58 ± 0.64 78.03 ± 1.40
6 Related works
Cooperative GNNs. While most GNNs operate by every node propagating information from or to its neighbors, there is at least one other work that breaks this paradigm. We reuse the terminology introduced in the CO-GNN$(\pi,\eta)$ model (Finkelshtein et al., 2024) and redesign the architecture via sheaf theory. The CO-GNN$(\pi,\eta)$ architecture is described by two cooperating GNNs: the GNN $\pi$ chooses the best action (isolate, listen, broadcast and standard) and the GNN $\eta$ updates node representations. Our architecture is described by two cooperating GNNs that learn the maps ${\mathbf{T}}_{i}$ and ${\mathbf{S}}_{i}$, which will simultaneously decide the action of the node $i$ and update it, for all $i\in V$. We highlight that our method avoids the Gumbel-softmax estimator Jang et al. (2017); Maddison et al. (2017), since we do not have an action network for predicting categorical actions for the nodes; instead, the learnable maps provide the actions by approaching or distancing to the null map.
Sheaf Neural Networks. Our model is inspired by the NSD model proposed by Bodnar et al. (2022) and influenced by the BuNN model (Bamberger et al., 2024), since we use only two restriction maps per node to avoid overfitting and increase scalability. Importantly, our model is closer to NSD than to BuNN: NSD originates from a discretization of the heat equation defined over cellular sheaves, leading to a standard message passing algorithm. Similarly, CSNN originates from a discretization of a modified heat equation that replaces the sheaf Laplacian by a composition of sheaf Laplacians, leading to a different message passing paradigm since nodes can choose not to listen or broadcast, providing the ability to mitigate over-squashing. In contrast, BuNNs uses the heat kernel directly, i.e., it replaces the sheaf Laplacian by $\exp(-tL_{\cal{B}})$, where $\cal{B}$ is a flat bundle. The diffusion behavior induced by this operator becomes more local or more global depending on the value of $t$, also mitigating over-squashing. Thus, we have two different sheaf-theoretic methods that alleviates over-squashing as a by-product of the model. Moreover, we contributed to sheaf neural networks providing sound sheaf Laplacians for directed graphs, which were crucial to model the actions of listening and broadcasting independently.
Quiver Laplacians. Sumray et al. (2024) defined sheaf Laplacians over quivers. In their context, quivers are directed graphs with self-loops. The in- and out-Laplacians we defined here are not particular cases of the Laplacians over quivers: they obtain a positive semi-definite matrix while our Laplacians can have complex eigenvalues with negative real parts.
7 Discussion
This work proposed Cooperative Sheaf Neural Networks, a novel SNN architecture that incorporates directionality in message-passing in order to increase its efficiency by learning sheaves with conformal maps, allowing nodes to choose the optimal behavior in terms of information propagation with respect to its neighbors. We provide theoretical insights on how CSNN can alleviate over-squashing due to this rewiring capacity, and test its effectiveness on node classification experiments for heterophilic graphs.
Limitations and Future works. There are works theoretically and experimentally studying the over-squashing issue in message-passing graph neural networks. Since SNNs are expansions of GNNs, discussing this phenomenon through the analysis of the Jacobian matrix of node features seems reasonable as a first step. However, further investigations must be made to understand to which extent the current tools to formally analyze over-squashing in GNNs are applicable to SNNs.
We must point out that the superior performance of our model comes at a price: CSNN has at least twice the computational cost of BuNN (Bamberger et al., 2024), since we learn two maps per node. This leads to scalability issues, a central problem of sheaf-based models, and alleviating this cost is a promising direction for future works. Finally, the sheaf Laplacians for directed graphs allow the use of sheaves to expand the relevance of directed graphs in neural networks (Rossi et al., 2024).
Broader impact. We do not foresee immediate negative societal impacts at this stage of the work.
Appendix A Experiment details
In this section, we provide the grid of hyperparameters used in the experiments. If the number of GNN layers is set to 0, we use an MLP with two layers to learn the restriction maps. Otherwise, we adopt a GraphSAGE architecture with the specified number of layers.
We also trained CO-GNN using the hyperparameter table from Finkelshtein et al. [2024], considering $\mu$ and $\Sigma$ as explicit hyperparameters instead of treating CO-GNN$(\mu,\mu)$ and CO-GNN(Σ,$(\Sigma,\Sigma)$ as separate model variants.
All datasets except for roman-empire were treated as undirected graphs. For the roman-empire dataset, we found that using the stored list of edges was preferable to doubling the edges, since the graphs from Platonov et al. [2023] are stored as "directed" lists where elements as (0,2) and (2,0) are regarded as equivalent, for example.
All experiments were conducted on a cluster equipped with NVIDIA V100, A100, and H100 GPUs. The choice of GPU depended on the availability at the time of the experiments. Each machine was provisioned with at least 80 GB of RAM.
Table 3: Hyperparameter configurations used across heterophilic benchmarks.
Parameter roman-empire, amazon-ratings minesweeper, tolokers, questions
sheaf dimension 3, 4, 5 3, 4, 5
# layers 2–5 2–5
# hidden channels 32, 64 32, 64
# of GNN layers 0–5 0–5
GNN dimension 32, 64 32, 64
dropout 0.2 0.2
input dropout 0.2 0.2
# epochs 2000 2000
activation GELU GELU
left weights true, false true, false
right weights true, false true, false
learning rate $0.02$ 0.002, 0.020.0020.020.002,\,0.020.002 , 0.02
[…]
We also present some statistics of the heterophilic benchmarks.
Table 4: Statistics of the heterophilous datasets
roman-empire amazon-ratings minesweeper tolokers questions
nodes 22662 24492 10000 11758 48921
edges 32927 93050 39402 519000 153540
avg degree 2.91 7.60 7.88 88.28 6.28
node features 300 300 7 10 301
classes 18 5 2 2 2
edge homophily 0.05 0.38 0.68 0.59 0.84
adjusted homophily -0.05 0.14 0.01 0.09 0.02
metric acc acc roc auc roc auc roc auc
Appendix B Robustness to over-squashing: the NeighborsMatch problem
In order to verify our theoretical results on the capacity of CSNN to alleviate over-squashing, we reproduce the NeighborsMatch problem proposed by Alon and Yahav [2020], using the same experimental setup. We recall that the datasets consist of binary trees of depth $r$, with the root node as the target, the leaves containing its possible labels, and the leaf with the same number of neighbors as the target node containing its true label. We provide the parameters used for this task below. The number of layers and hidden channels are shared between all models in the experiment.
Table 5: Hyperparameter configuration used for NeighborsMatch.
Parameter NeighborsMatch
sheaf dimension 2
# layers $r+1$
# hidden channels 32
# of GNN layers $r+1$
GNN dimension 32
dropout 0.0
input dropout 0.0
activation Id
left weights true
right weights true
layer norm true